OpenStudy (anonymous):

time is running out when mike is taking a multiple choice test so he guesses on the last 5 questions. Each question has 4 possible choices. what is the probability that mike gets exactly 3 of these 5 questions correct? 1)100% 2) 60% 3)10.4% 4)8.8%

OpenStudy (anonymous):

Why did he guess?

OpenStudy (anonymous):

Why did mike take so long on the rest of his test?

jimthompson5910 (jim_thompson5910):

use the binomial distribution formula P(X = k) = (n C k)*(p)^(k)*(1-p)^(n-k) In this case n = 5 p = 0.25 k = 3 P(X = k) = (n C k)*(p)^(k)*(1-p)^(n-k) P(X = 3) = (5 C 3)*(0.25)^(3)*(1-0.25)^(5-3) P(X = 3) = (10)*(0.25)^(3)*(1-0.25)^(5-3) I'll let you finish up

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

np